OpenStudy (dls):

Circles are drawn passing through the points(2,0) which make intercept of 5 unit length on X-axis.If their centres lie in 1st quad,eq of circle=?

OpenStudy (dls):

\[\LARGE 5=2 \sqrt{g^2-c}\] \[\LARGE \frac{25}{4}=(g^2-c)\]

Parth (parthkohli):

:nothingtodohere:

OpenStudy (dls):

lol

OpenStudy (dls):

@rajathsbhat

OpenStudy (anonymous):

LOL

OpenStudy (dls):

:o

OpenStudy (dls):

i freaking love that meme :P

OpenStudy (anonymous):

are we missing a point here?

OpenStudy (anonymous):

there must be a second point i assume...

OpenStudy (dls):

 arent there 2 cases possible? and no..

OpenStudy (anonymous):

i guess i'm not understanding the q here.. :\

OpenStudy (dls):

\[\LARGE x^2+y^2+2gx+2fy+g^2-\frac{25}{4}=0\]

OpenStudy (dls):

\[4+4g+g^2-\frac{25}{4}\] on subbing (2,0)

OpenStudy (dls):

arre hogaya sorry :o